OpenStudy (anonymous):

Complete the square and find the maximum value of f(x)=1/(x^2+2x+2).

OpenStudy (anonymous):

The answer is 1 but idk how to solve it

OpenStudy (anonymous):

i assume it means complete the square in the denominator it is not necessary to do that to answer this question, but we can do it if you like there is a much easier way

OpenStudy (anonymous):

\[f(x)=\frac{1}{x^2+2x+2}=\frac{1}{x^2+2x+1+1}=\frac{1}{(x+1)^2+1}\]

OpenStudy (anonymous):

now the the square has been completed, we can reason that the smallest the denominator can be is \(1\) which it is if \(x=-1\) then since the smallest the denominator can be is 1, the largest the entire function can be is \[f(-1)=\frac{1}{1}=1\]

OpenStudy (anonymous):

we could have done the same work by minimizing \[x^2+2x+2\] which has a minimum at the vertex the first coordinate of the vertex is \(-\frac{b}{2a}=-\frac{2}{2}=-1\) and the second coordinate of the vertex is \((-1)^2+2(-1)+2=1\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw